# Large Language Models for Blockchain Security: A Systematic Literature Review

### v3

@@ 3.6 Miscellaneous
As we display in Table 7, LLM is also used in blockchain security fields, involving smart contract compilers [45], zero-knowledge proofs [46], model training [20, 47], NFT generation [48]. We will introduce their applications in detail in future work.
@@ 4.1 LLM4Fuzz
Figure 4: The architecture of LLM4FUZZ.
As we depicted in Fig.4, LLM4FUZZ [35] emerges as an innovative technique in the cybersecurity landscape, specifically in the niche of smart contract security within blockchain networks. It intricately combines the prowess of Large Language Models (LLMs) with fuzz testing methodologies to proactively unearth vulnerabilities that could potentially compromise the integrity of smart contracts.
@@ 4.2 SMARTINV
Proposed with the intention of enhancing the reliability and security of blockchain smart contracts, SMARTINV [24] represents a significant breakthrough in the field. Its primary function is to infer invariants within smart contracts which can be integral in automating the process of identifying elusive bugs that typically elude conventional machine-auditing methods. Fig.5 display the architecture of SMARTINV.
@@ 4.3 BLOCKGPT
As we showed in Fig.6, BLOCKGPT [21] serves as a paradigm shift in the domain of blockchain security, acting as a state-of-the-art Intrusion Detection System (IDS) specifically engineered to counteract and identify potentially malicious transactions within blockchain networks. The system is underpinned by a highly sophisticated large language model that has been meticulously trained with a significant corpus of transactional data from the Ethereum blockchain, one of the most widely utilized platforms in the industry.
@@
5 Future Direction and Challenge of LLM4BS
@@
Regulatory and Compliance Challenges: The shifting sands of regulatory frameworks demand not only compliance but a proactive engagement with regulatory bodies by scholars and practitioners in the LLM4BS field [5, 113]. This relationship is reciprocal; as regulatory agencies develop deeper understandings of the implications of integrating AI in blockchain, it is incumbent upon the actors within this space to advocate for regulations that encourage innovation while maintaining robust security measures. The dynamic interplay between cutting-edge technology and regulation is a delicate balance to strike, fostering a stable yet flexible platform for growth and adaptation in blockchain security solutions.
@@
Data Quality and Access: At the heart of robust LLM4BS deployments lies the foundational element of data—its caliber, its scope, and the accessibility afforded to it. Herein lies the challenge: constructing and maintaining databases that are not only comprehensive and representative but are also curated with an eye towards enhancing the efficacy of Large Language Models in detecting anomalies and reinforcing security parameters in blockchain transactions. The task extends to crafting protocols that ensure data integrity and sourcing that conforms to ethical standards, thereby upholding the sanctity and reliability of these AI systems.
@@ 6 Conclusion
In summing up this review on the incorporation of Large Language Models (LLMs) into blockchain security, we have traversed through the technological advancements and intricate challenges presented by this combination of LLM4BS. The potential of LLMs to augment security protocols in the blockchain is clear, offering innovative solutions for smart contracts, detection of abnormal transactions, and development of the cryptocurrency community. Yet, this potential comes with the necessity for vigilance regarding scalability, privacy, advancing cyber threats, and ethical implications of AI. As we look ahead, the success of LLMs in blockchain security depends not only on continuous technological refinement but also on ethical practices, regulatory alignment, and informed community engagement. The integration of LLM into blockchain security heralds a transformative era that demands a collaborative approach, balancing innovation with prudent oversight to forge a resilient and equitable security future.

### v4

@@ 3.6 Miscellaneous
As displayed in Table 7, LLM is also used in other blockchain security fields, involving smart contract compilers [57], zero-knowledge proofs [58], model training [20, 59], NFT generation [60]. We will introduce their applications in detail in the future.
@@ 4.1 LLM4Fuzz
As depicted in Fig.4, LLM4FUZZ [44] emerges as an innovative technique in the cybersecurity landscape, specifically in the niche of smart contract security within blockchain networks. It intricately combines the prowess of Large Language Models (LLMs) with fuzz testing methodologies to proactively unearth vulnerabilities that could potentially compromise the integrity of smart contracts.
@@ 4.2 SMARTINV
Proposed with the intention of enhancing the reliability and security of blockchain smart contracts, SMARTINV [29] represents a significant breakthrough in the field. Its primary function is to infer invariants within smart contracts, which can be integral in automating the process of identifying elusive bugs that typically elude conventional machine-auditing methods. Fig.5 displays the architecture of SMARTINV.
@@ 4.3 BLOCKGPT
As shown in Fig.6, BLOCKGPT [26] serves as a paradigm shift in the domain of blockchain security, acting as a state-of-the-art Intrusion Detection System (IDS) specifically engineered to counteract and identify potentially malicious transactions within blockchain networks. The IDS is underpinned by a highly sophisticated large language model that has been meticulously trained with a significant corpus of transactional data from the Ethereum blockchain, one of the most widely utilized platforms in the industry.
@@
5 Future Direction and Challenge of LLM4BS tasks
@@
Regulatory and Compliance Challenges: The shifting sands of regulatory frameworks demand not only compliance but a proactive engagement with regulatory bodies by scholars and practitioners in the LLM4BS field [5, 125]. This relationship is reciprocal; as regulatory agencies develop a deeper understanding of the implications of integrating AI in blockchain, it is incumbent upon the actors within this space to advocate for regulations that encourage innovation while maintaining robust security measures. The dynamic interplay between cutting-edge technology and regulation is a delicate balance to strike, fostering a stable yet flexible platform for growth and adaptation in blockchain security solutions.
@@
Data Quality and Access: At the heart of robust LLM4BS deployments lies the foundational element of data—its caliber, its scope, and the accessibility afforded to it. Herein lies the challenge: constructing and maintaining databases that are not only comprehensive and representative but are also curated with an eye toward enhancing the efficacy of Large Language Models in detecting anomalies and reinforcing security parameters in blockchain transactions. The task extends to crafting protocols that ensure data integrity and sourcing that conforms to ethical standards, thereby upholding the sanctity and reliability of these AI systems.
@@ 6 Conclusion
In conclusion, our review of the integration of Large Language Models (LLMs) into blockchain security highlights the technological advancements and intricate challenges presented by this combination of LLM4BS. The potential of LLMs to enhance security protocols in the blockchain is evident, offering innovative solutions for smart contracts, abnormal transaction detection, and cryptocurrency community development. However, realizing this potential requires vigilance regarding scalability, privacy, evolving cyber threats, and the ethical implications of AI. The success of LLMs in blockchain security hinges not only on continuous technological refinement but also on ethical practices, regulatory alignment, and informed community engagement. The integration of LLMs into blockchain security marks a transformative era that necessitates a collaborative approach, balancing innovation with prudent oversight to forge a resilient and equitable security future.